OpenStudy (anonymous):

-3<(2m+4/2)< or equal to 10 solve

OpenStudy (anonymous):

Do you mean -3<(2m+4)/2< or equal to 10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-3<(2m+4)/2 =< 10 -3<m+2 =< 10 need to break up the problem into two different sets. -3<m+2 -5 < m and m+2 =< 10 m=< 8 then m is said to be -5 < m and m=< 8 or -5 < m=< 8

OpenStudy (anonymous):

i can kiss you

OpenStudy (anonymous):

tanks